OpenStudy (toxicsugar22):

ater (3110g ) is heated until it just begins to boil. If the water absorbs 5.01×105J of heat in the process, what was the initial temperature of the water?

OpenStudy (toxicsugar22):

do you get it

OpenStudy (sinbearciante):

Using dimensional analysis, I came up with 334.516K as the initial temperature. I am going to try and check it, though, with an eq.

OpenStudy (sinbearciante):

I also got the same answer when using the heat capacity eq, C=dQ/dT. Plugging the knowns in, I found that the dT was 119.685 K*kg. I then divided that by the kg of water present to find the change in temp and then subtracted that from 373K.

OpenStudy (sinbearciante):

I feel fairly confident in my answer, but hopefully somebody can provide an answer that uses a more formal eq.

OpenStudy (toxicsugar22):

ohh okay so answe is 334.516K

OpenStudy (sinbearciante):

I believe so

OpenStudy (toxicsugar22):

IT SAYS TRY USIND A DIFFERENT TEMP SCALE

OpenStudy (anonymous):

my sln: q = m*Cs*(T2 - T1) (in our case T2 = 373K boiling T) 5.05*10^5J = 3110*4.18*(373-T1) from there T1 = 334.2K 334.2K - 273 ~ 61 Celsius

OpenStudy (toxicsugar22):

61 IS ANSWER

OpenStudy (toxicsugar22):

IT SAYS IT WANT USING DIMENSIONS OF TEMP

OpenStudy (anonymous):

I wrote that..334K is 61Celsius degree... 334K - 273 = 61 C

OpenStudy (anonymous):

K - 273 = C

OpenStudy (toxicsugar22):

334